In 1895, the first a sporting event was held. The winner's prize money was $190. In 2007, the winner's check was $1,162,000. (Do not round your intermediate calculations.)

Required:
(a) What was the percentage increase per year in the winner's check over this period?
(Click to select)  8.20  8.15  8.10  8.00  -23.22
(b) If the winner's prize increases at the same rate, what will it be in 2040?
   (Click to select)  190.00  14,709,326.81  15,635,420.86  15,398,732.66  15,165,520.45
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.

1,162,000=190*(1+r/100)

(1,162,000/190)^(1/112)=(1+r/100)

(1+r/100)=1.081

r=1.081-1

=8.1%(Approx).

b.A=1,162,000*(1.081)^33

=$15165520.45(Approx).